OpenStudy (anonymous):

Is anyone good at physical science speed, velocity, and acceleration practice problems

OpenStudy (anonymous):

That's what Calculus and Physics is all about. Lay it on me! (I hope I can help o.0)

OpenStudy (anonymous):

1. what is the velocity of a plane that traveled 3,000 miles from New york to California 5.0 hours? Known Values: Unknown Values: Equation to solve: Substitutions: Solution:

OpenStudy (stormfire1):

Average Velocity is displacement/time so it's 3000mi/5hrs = 600 mph

OpenStudy (anonymous):

3,000 miles per 5.0 hours Simplify that down by dividing 3,000 by 5 to get an even "mph" value. \[{3000\over5}=600\] 600 miles per hour (mph)

OpenStudy (anonymous):

what are the known values

OpenStudy (stormfire1):

3000 miles and 5 hours

OpenStudy (stormfire1):

and the equation is V = Displacement/Time...and you have displacement and time

OpenStudy (stormfire1):

more correctly: V(avg) = Displacement/Time

OpenStudy (anonymous):

what are the unknown values???

OpenStudy (stormfire1):

Velocity? heh

OpenStudy (anonymous):

does the velocity = d/v go for the unknown value

OpenStudy (stormfire1):

I guess you could say that. You're trying to solve for v (the unknown) using the known values of d and t

OpenStudy (anonymous):

what's the equation to solve it

OpenStudy (stormfire1):

v = d/t

OpenStudy (anonymous):

so put that equation for both unknown and he equation to solve

OpenStudy (stormfire1):

I'm not sure what's unclear here... you're solving for V and you know that V = distance divided by time...both values which you already have. From there, it's just a simple division problem

OpenStudy (anonymous):

the equation to solve I mean

OpenStudy (anonymous):

what are the substitutions

OpenStudy (stormfire1):

v = 3000/5 shows your substitutions for the known values...you just need to perform the division to get 600 which = the average velocity